# Continual Learning in Medical Imaging Analysis: A Comprehensive Review of Recent Advancements and Future Prospects

### v1

@@ 3.3 Task incremental scenarios
Task Incremental Scenario (TIS) comes into the picture when we have a multi-task problem and a single adaptive model is desired. Each task is considered an episode. Thus each episode has disjoint label space. However, there is ambiguity in the literature with task incremental scenarios as the other scenarios (class and domain) are frequently referred to as task incremental scenarios by different research communities. For example, Ravishankar et al. (2019) create the first episode having two chest x-ray views (2ch vs. PLAX) and then the next episode with two other views (4ch vs. PSAX) as a task incremental scenario which can also be a class incremental scenario. Further another example in this line can be found in the work by Baweja et al. (2018). Authors create a task incremental scenario with two episodes for brain MRI segmentation application where the $1^{st}$ episode was multi-class segmentation of cerebrospinal fluid, grey matter, and white matter, and the $2nd$ episode was the segmentation of white matter lesions. If we have multiple tasks that are very different then it would certainly be the task incremental; however, if the tasks are more close then it is merely a design choice whether to treat them as class incremental or task incremental. Another work Kaustaban et al. (2022) explores domain incremental scenario as organ shift for tumor classification problem where the first episode is a colon cancer dataset with 9 classes (CRC dataset) and the next episode is a breast cancer dataset with 2 classes (PatchCam dataset). Similar is the case with Sadafi et al. (2023) where WBC classification from three datasets, each having a different number of classes is regarded as domain incremental which can also be regarded as task-incremental. Derakhshani et al. (2022) curate class incremental and task incremental scenarios from the same dataset taken from MedMNIST Yang et al. (2023a) where a dataset is divided into non-overlapping classes leading to 4 episodes. If the model is evaluated only on the learnt class in the specific episode then they regard it as task incremental else if evaluated on cumulative classes from all seen episodes then class incremental scenario.
Kaustaban et al. (2022) say that any of the scenarios within data, class, and domain incremental can be treated as task incremental, provided that each incoming data stream (episode) is treated as a distinct task. For a comprehensive overview, we have tabulated all the works along with their claimed task incremental scenario in Table 4.
Further, it is essential to consistently provide prior information about which specific task (referred to as task identity, i.e., task ID) the test data pertains to, and predictions are made accordingly. However, if task IDs are not provided then this category is termed as a task-free scenario which is a more challenging one.
@@
sMRI based Axial and sagittal brain segmentation: datasets from 4 sites: HCP (Van Essen et al. 2013), NKI (Nooner et al. 2012), Buckner (Biswal et al. 2010), WU120 (Power et al. 2017) in different order
Karani et al. (2018)
cross-scanners, hospitals, or acquisition protocols (4 Ep.)
@@
MR brain segmentation on various combination using 5 datasets from different source: Human Connectome Project (Van Essen et al. 2013), Alzheimer’s Disease Neuroimaging Initiative (ADNT ), Autism Brain Imaging Data Exchange (Di Martino et al. 2014) and Information eXtraction from Images (IXI )
van Garderen et al. (2019)
Low vs. high-grade (2 Ep.)
@@
MR data based glioma segmentation: 2018 BraTS Challenge (low and high-grade), an in-house dataset with non-enhancing low-grade
Venkataramani et al. (2019)
varying in disease type, intensity patterns (2) & contrast
@@
X-ray lung segmentation: Montgomery (Jaeger et al. 2014) as source, JSRT (Shiraishi et al. 2000) and Pneumoconiosis (private) as two target domains
Ravishankar et al. (2019)
Inter-subject variability (2 Ep.)
@@
(a) Chest x-ray view classification (4ch vs. PLAX) from adult to pediatric as two domains (b) chest x-ray view classification (2ch vs. PSAX) from adult to pediatric as two domains
Lenga et al. (2020)
Cross-sites (2 Ep.)
@@
Chest x-ray classification on datasets from 2 source: ChestX-ray14 dataset from NIH Clinical Center and MIMIC-CXR dataset from the Beth Israel Deaconess Medical Center
Özgün et al. (2020)
Different age ranges, MRI field strengths & presence of pathologies or motion artifacts (total 4 Ep.)
@@
Brain MRI segmentation datasets: 4 Ep. were created from 3 datasets: CANDI (Kennedy et al. 2012) (1 Ep.), ADNI (Jack Jr et al. 2008) (1 Ep.), and MALC (Asman and Landman 2013) (2 Ep. based on age)
Hofmanninger et al. (2020a)
Total 3 Ep. (a) different scanner parmeter (b) target-shift
@@
Chest CT synthetic classification task: (a) first to second episode have change in scanner parameter (Hofmanninger et al. 2020b) (b) target shift is introduced intentionally in the third episode by imprinting a synthetic target structure in the form of a cat on random locations, rotations and varying scale in 50% of the samples.
Morgado et al. (2021)
Appearance and body-parts (2 Ep.)
@@
Dermatological imaging modality classification (full-body, anatomic, macroscopic, and dermoscopic classes): For each of the class author considered disjoint images in the two episodes, for example, in the second episode, legs and arms images were considered for full-body class, images containing hands or feet only considered for anatomic, images containing face region for macroscopic class, and pink colored images for dermoscopic class.
Srivastava et al. (2021)
Cross-sites (3 Ep.)
@@
Chest x-ray classification: NIH Chest-X-rays14, PadChest, and CheXpert
Memmel et al. (2021)
Multi-scanner (2 Ep.)
@@
MRI-based Hippocampal Segmentation: 3 datasets as 2018 Medical Segmentation Decathlon challenge (Simpson et al. ), Scientific Data (Kulaga-Yoskovitz et al. 2015), and Alzheimer’s Disease Neuroimaging Initiative (Boccardi et al. 2015): 2 datasets together in 1st episode and then 3rd dataset in 2nd episode, different such combinations were experimented
Perkonigg et al. (2021)
Multi-scanners (3 Ep.)
@@
Brain age estimation: IXI(IXI ) (Philips Gyroscan Intera 1.5T, Philips Intera 3.0T scanner), OASIS3 (LaMontagne et al. 2019) (Siemens TrioTim 3.0T scanner)
Bayasi et al. (2021)
Multi-site (6 Ep.)
@@
Skin lesion image classification: HAM10000 (Tschandl et al. 2018), Dermofit (Ballerini et al. 2013), Derm7pt (Kawahara et al. 2018), MSK (Codella et al. 2018), PH2 (Mendonça et al. 2013), UDA (Codella et al. 2018)
Zhang et al. (2021)
Cross-sites (4-6 Ep.)
@@
(a) Prostate segmentation: MRI datasets across 6 sites: RUNMC (Bloch et al. 2015), BMC (Bloch et al. 2015), HCRUDB (Lemaître et al. 2015), UCL (Litjens et al. 2014), BIDMC (Litjens et al. 2014), HK (Litjens et al. 2014), (b) optic cup and disc segmentation: public multi-site fundus image datasets from 4 sources (Sivaswamy et al. 2015; Fumero et al. 2011; Orlando et al. 2020)
Perkonigg et al. (2022)
Multi-scanner (4 Ep.)
@@
(a) CMR based Cardiac segmentation (4 Ep.): Cardiac (Campello et al. 2021), (b) CT based Lung nodule detection (4 Ep.): LIDC (Setio et al. 2017), + LNDb challenge (Pedrosa et al. 2019), (c) MRI based Brain Age Estimation segmentation (4 Ep.): IXI(IXI ) + OASIS-3 (LaMontagne et al. 2019)
Derakhshani et al. (2022)
Across organs or modality (4 Ep.)
@@
disease classification: 4 epi using 4 dataset ordered as: BloodMNIST, OrganaMNIST, PathMNIST, TissueMNIST (MedMNIST repository)
Karthik et al. (2022)
Cross-center (8 Ep.)
@@
Sclerosis lesions segmentation from brain MRI datasets described in Kerbrat et al. (Kerbrat et al. 2020)
González et al. (2022)
(a) Cross-domain (5 Ep.) (b) manual image contrast change (5 Ep.)
@@
MRI hippocampus segmentation: (a) HarP (Boccardi et al. 2015), Dryad (Kulaga-Yoskovitz et al. 2015), Decathlon (Simpson et al. 2019) (b) MRI hippocampus segmentation: data transformation applied using TorchIO library on Decathlon (Simpson et al. 2019) as intensity rescaling, affine transformations, rotation, translation
Ranem et al. (2022)
Cross-site (3 Ep.)
@@
MRI based binary hippocampus segmentation: Decathlon (Antonelli et al. 2022), Drayd (Denovellis et al. 2021), HarP (Boccardi et al. 2015)
Kaustaban et al. (2022)
(a) Multi-organ (2 Ep.) (b) scanning protocol (5 Ep.)
@@
(a) Tumor classification: CRC (a colon cancer dataset with 9 classes), PatchCam (a breast cancer dataset with 2 classes) (b) 5 domain scenario were simulated by changing H&E composition in CRC dataset
Shu et al. (2022)
(a) Imaging protocol (2 Ep.) (b) multi-source (3 epi)
@@
(a) Low and high-quality fundus images (retinal images) from EyeQ dataset as domain shift condition (Fu et al. 2019), (b) fundus disease classification (AMD, DR, glaucoma, myopia, and normal classes): ODIR (ODIR ), R&R (RIADD (RIADD ) + REFUGE (REFUGE )), and iSee (Fang et al. 2020)
Li et al. (2022)
Multi-scanner vendors (4 Ep.)
@@
CMR based Cardiac segmentation of left ventricle, right ventricle, and left ventricle myocardium: 4 scanner vendors (Siemens, Philips, General Electric, Cannon) from M&Ms (Campello et al. 2021) dataset
Sadafi et al. (2023)
Cross-sites (3 Epi)
WBC classification datasets: Matek-19 (Matek et al. 2019), INT-20, and Acevedo-20 (Acevedo et al. 2020)
Bera et al. (2023)
(a) Cross-center (4 Ep.) (b) cross-center (2 Ep.)
@@
(a) Binary prostate segmentation: Prostate158 (Adams et al. 2022) $\rightarrow$ NCI-ISBI (NCI ) $\rightarrow$ Promise12 (Litjens et al. 2014) $\rightarrow$ Decathlon (Antonelli et al. 2022) (b) binary hippocampus segmentation: Drayd (Denovellis et al. 2021) $\rightarrow$ HarP (Boccardi et al. 2015))
Zhu et al. (2023)
(a) Cross-site (6 Ep.) (b) cross-site & cross-modality (2 Ep.) (c) same-site & cross-modality (2 Ep.)
@@
(a) Binary prostate segmentation from T2-weighted MRI scans collected from 6 sites (12–30 scans/site) (Liu et al. 2020; Bloch et al. 2015; Lemaître et al. 2015) (b) multi-class (liver, left and right kidneys, and spleen) abdominal segmentation between 30 CT and 20 MRI T2-SPIR scans (c) muscle segmentation of 13 lower-leg muscles and bones between 30 MRI T1 and 30 mDixon scans
Bándi et al. (2023)
Cross-organ (3 Ep.)
@@
Histopathology data based cancer detection from breast (CAMELYON16, CAMELYON17), colon (private), and head-neck (private) datasets
@@ 3.4 Domain incremental scenarios
Domain Incremental Scenario (DIS) is the most popular and frequently observed category of CL scenarios for medical applications. This scenario aligns with the idea of continual learning in a changing environment where the datasets from different domains (e.g., research sites, hospitals, imaging modality, image acquisition protocol, etc.) are incrementally encountered over time which thus involve covariance shift-induced discrepancies in data.
@@
Even if we have a single task and a fixed number of classes, these shifts can pose a challenge for the static deep model used in clinical settings because these shifts can quickly make existing models outdated and less effective. Then a CL model aims to continuously update itself by incorporating new data streams that come from distributions that have shifted over time.
There have been plenty of attempts to develop CL models and evaluate different kinds of domain incremental scenarios. All the domain incremental scenarios are not equi-hard; some works use a simpler level of domain shift whereas some with severe domain shift. To provide a comparative view, we tabulate all the domain incremental scenarios considered in literature through Table 5.
@@
Liu et al. (2023a)
Brain tumor segmentation (3 Ep.)
CIS with domain shift conditions: incrementally learn tumor core (BraTS2013 (Menze et al. 2014) dataset), the enhancing tumor (TCIA (Clark et al. 2013) dataset), and edema (CBICA (Bakas et al. 2018) dataset) structures via three datasets, each following different data distributions
Sadafi et al. (2023)
WBC classification (3 Ep.)
DIS (multi-site) + CIS using 3 WBC classification datasets having different number of classes: CIS on Matek-19 (Matek et al. 2019) then shift to INT-20 dataset and use it in CIS manner then shift to Acevedo-20 (Acevedo et al. 2020) dataset and use in CIS manner
Yang et al. (2023b)
MedMNIST (Yang et al. 2023a) based disease classification (4 Ep.)
CIS+DIS: 3 datasets (PathMNIST, DermaMNIST, and OrganAMNIST) as source domain and then 3 more datasets (RetinaMNIST, BreastMNIST, BloodMNIST) as domain shift but classes are added in class incremental fashion with 1 class at a time
Wang et al. (2023b)
CIS (a) segmentation for endoscopy (2-3 Ep.) (b) surgical instrument segmentation (2 Ep.)
Incrementally adding segmentation structure from EDD2020 (Ali et al. 2021, 2020) dataset which includes 5 classes as Barrett’s esophagus, cancer, high-grade dysplasia, polyp, and suspicious. Initially, 3-4 classes are in 1st episode, and then 2 or 1 classes are incrementally added in subsequent episodes. (b) EndoVis18 (Allan et al. 2020) and EndoVis17 (Allan et al. 2019) datasets as two episodes having some overlapping classes
Liu et al. (2023a)
MRI-based brain tumor segmentation (4 Ep.)
CIS+DIS Incrementally adding more structures to segmentation task from dataset acquired from different vendors/centers

### v3

@@ 3.3 Task incremental scenarios
Task Incremental Scenario (TIS) comes into the picture when we have a multi-task problem and a single adaptive model is desired. Each task is considered an episode. Thus each episode has disjoint label space. However, there is ambiguity in the literature with task incremental scenarios as the other scenarios (class and domain) are frequently referred to as task incremental scenarios by different research communities. For example, Ravishankar et al. (2019) create the first episode having 2 chest x-ray views (2ch vs. PLAX) and then the next episode with 2 other views (4ch vs. PSAX) as a task incremental scenario, which can also be a class incremental scenario. Further, another example in this line can be found in the work by Baweja et al. (2018). Here, the authors create a task incremental scenario with 2 episodes for brain MRI segmentation application where the $1^{st}$ episode was multi-class segmentation of cerebrospinal fluid, grey matter, and white matter, and the $2nd$ episode was the segmentation of white matter lesions. If we have multiple tasks that are very different, then it would certainly be the task incremental; however, if the tasks are more close then it is merely a design choice whether to treat them as class incremental or task incremental. Another work where Kaustaban et al. (2022) explore domain incremental scenario as organ shift for tumor classification problem where the first episode is a colon cancer dataset with nine classes (CRC dataset), and the next episode is a breast cancer dataset with 2 classes (PatchCam dataset). Similar is the case with Sadafi et al. (2023) where White Blood Cells (WBC) classification from 3 datasets, each having a different number of classes, is regarded as domain incremental, which can also be regarded as task-incremental. Derakhshani et al. (2022) curate class incremental and task incremental scenarios from the same dataset taken from MedMNIST Yang et al. (2023a) where a dataset is divided into non-overlapping classes leading to 4 episodes. If the model is evaluated only on the learnt class in the specific episode, then they regard it as task incremental; otherwise, if evaluated on cumulative classes from all seen episodes, then class incremental scenario.
Kaustaban et al. (2022) say that any of the scenarios among instance, class, and domain incremental can be treated as task incremental, provided that each incoming data stream (episode) is treated as a distinct task. For a comprehensive overview, we have tabulated all the works along with their claimed task incremental scenario in Table 4.
Further, it is essential to consistently provide prior information about which specific task (referred to as task identity, i.e., task ID) the test data pertains to, and predictions are made accordingly. However, if task IDs are not provided, then this category is termed as a task-free scenario which is a more challenging one.
@@
McClure et al. (2018) Cross-site (4 Ep.) sMRI based Axial and sagittal brain segmentation: datasets from 4 sites: HCP (Van Essen et al. 2013), NKI (Nooner et al. 2012), Buckner (Biswal et al. 2010), WU120 (Power et al. 2017) in different order
@@
Karani et al. (2018) cross-scanners, hospitals, or acquisition protocols (4 Ep.) MR brain segmentation on various combination using 5 datasets from different source: Human Connectome Project (Van Essen et al. 2013), Alzheimer’s Disease Neuroimaging Initiative (ADNT ), Autism Brain Imaging Data Exchange (Di Martino et al. 2014) and Information eXtraction from Images (IXI )
@@
van Garderen et al. (2019) Low vs. high-grade (2 Ep.) MR data based glioma segmentation: 2018 BraTS Challenge (low and high-grade), an in-house dataset with non-enhancing low-grade
@@
Venkataramani et al. (2019) varying in disease type, intensity patterns (2) & contrast X-ray lung segmentation: Montgomery (Jaeger et al. 2014) as source, JSRT (Shiraishi et al. 2000) and Pneumoconiosis (private) as two target domains
@@
Ravishankar et al. (2019) Inter-subject variability (2 Ep.) (a) Chest x-ray view classification (4ch vs. PLAX) from adult to pediatric as two domains (b) chest x-ray view classification (2ch vs. PSAX) from adult to pediatric as two domains
@@
Lenga et al. (2020) Cross-sites (2 Ep.) Chest x-ray classification on datasets from 2 source: ChestX-ray14 dataset from NIH Clinical Center and MIMIC-CXR dataset from the Beth Israel Deaconess Medical Center
@@
Özgün et al. (2020) Different age ranges, MRI field strengths & presence of pathologies or motion artifacts (total 4 Ep.) Brain MRI segmentation datasets: 4 Ep. were created from 3 datasets: CANDI (Kennedy et al. 2012) (1 Ep.), ADNI (Jack Jr et al. 2008) (1 Ep.), and MALC (Asman and Landman 2013) (2 Ep. based on age)
@@
Hofmanninger et al. (2020a) Total 3 Ep. (a) different scanner parmeter (b) target-shift Chest CT synthetic classification task: (a) first to second episode have change in scanner parameter (Hofmanninger et al. 2020b) (b) target shift is introduced intentionally in the third episode by imprinting a synthetic target structure in the form of a cat on random locations, rotations and varying scale in 50% of the samples.
@@
Morgado et al. (2021) Appearance and body-parts (2 Ep.) Dermatological imaging modality classification (full-body, anatomic, macroscopic, and dermoscopic classes): For each of the class author considered disjoint images in the two episodes, for example, in the second episode, legs and arms images were considered for full-body class, images containing hands or feet only considered for anatomic, images containing face region for macroscopic class, and pink colored images for dermoscopic class.
@@
Srivastava et al. (2021) Cross-sites (3 Ep.) Chest x-ray classification: NIH Chest-X-rays14, PadChest, and CheXpert
@@
Memmel et al. (2021) Multi-scanner (2 Ep.) MRI-based Hippocampal Segmentation: 3 datasets as 2018 Medical Segmentation Decathlon challenge (Simpson et al. ), Scientific Data (Kulaga-Yoskovitz et al. 2015), and Alzheimer’s Disease Neuroimaging Initiative (Boccardi et al. 2015): 2 datasets together in 1st episode and then 3rd dataset in 2nd episode, different such combinations were experimented
@@
Perkonigg et al. (2021) Multi-scanners (3 Ep.) Brain age estimation: IXI(IXI ) (Philips Gyroscan Intera 1.5T, Philips Intera 3.0T scanner), OASIS3 (LaMontagne et al. 2019) (Siemens TrioTim 3.0T scanner)
@@
Bayasi et al. (2021) Multi-site (6 Ep.) Skin lesion image classification: HAM10000 (Tschandl et al. 2018), Dermofit (Ballerini et al. 2013), Derm7pt (Kawahara et al. 2018), MSK (Codella et al. 2018), PH2 (Mendonça et al. 2013), UDA (Codella et al. 2018)
@@
Zhang et al. (2021) Cross-sites (4-6 Ep.) (a) Prostate segmentation: MRI datasets across 6 sites: RUNMC (Bloch et al. 2015), BMC (Bloch et al. 2015), HCRUDB (Lemaître et al. 2015), UCL (Litjens et al. 2014), BIDMC (Litjens et al. 2014), HK (Litjens et al. 2014), (b) optic cup and disc segmentation: public multi-site fundus image datasets from 4 sources (Sivaswamy et al. 2015; Fumero et al. 2011; Orlando et al. 2020)
@@
Perkonigg et al. (2022) Multi-scanner (4 Ep.) (a) CMR based Cardiac segmentation (4 Ep.): Cardiac (Campello et al. 2021), (b) CT based Lung nodule detection (4 Ep.): LIDC (Setio et al. 2017), + LNDb challenge (Pedrosa et al. 2019), (c) MRI based Brain Age Estimation segmentation (4 Ep.): IXI(IXI ) + OASIS-3 (LaMontagne et al. 2019)
@@
Derakhshani et al. (2022) Across organs or modality (4 Ep.) disease classification: 4 epi using 4 dataset ordered as: BloodMNIST, OrganaMNIST, PathMNIST, TissueMNIST (MedMNIST repository)
@@
Karthik et al. (2022) Cross-center (8 Ep.) Sclerosis lesions segmentation from brain MRI datasets described in Kerbrat et al. (Kerbrat et al. 2020)
@@
González et al. (2022) (a) Cross-domain (5 Ep.) (b) manual image contrast change (5 Ep.) MRI hippocampus segmentation: (a) HarP (Boccardi et al. 2015), Dryad (Kulaga-Yoskovitz et al. 2015), Decathlon (Simpson et al. 2019) (b) MRI hippocampus segmentation: data transformation applied using TorchIO library on Decathlon (Simpson et al. 2019) as intensity rescaling, affine transformations, rotation, translation
@@
Ranem et al. (2022) Cross-site (3 Ep.) MRI based binary hippocampus segmentation: Decathlon (Antonelli et al. 2022), Drayd (Denovellis et al. 2021), HarP (Boccardi et al. 2015)
@@
Kaustaban et al. (2022) (a) Multi-organ (2 Ep.) (b) scanning protocol (5 Ep.) (a) Tumor classification: CRC (a colon cancer dataset with 9 classes), PatchCam (a breast cancer dataset with 2 classes) (b) 5 domain scenario were simulated by changing H&E composition in CRC dataset
@@
Shu et al. (2022) (a) Imaging protocol (2 Ep.) (b) multi-source (3 epi) (a) Low and high-quality fundus images (retinal images) from EyeQ dataset as domain shift condition (Fu et al. 2019), (b) fundus disease classification (AMD, DR, glaucoma, myopia, and normal classes): ODIR (ODIR ), R&R (RIADD (RIADD ) + REFUGE (REFUGE )), and iSee (Fang et al. 2020)
@@
Li et al. (2022) Multi-scanner vendors (4 Ep.) CMR based Cardiac segmentation of left ventricle, right ventricle, and left ventricle myocardium: 4 scanner vendors (Siemens, Philips, General Electric, Cannon) from M&Ms (Campello et al. 2021) dataset
Sadafi et al. (2023) Cross-sites (3 Epi) WBC classification datasets: Matek-19 (Matek et al. 2019), INT-20, and Acevedo-20 (Acevedo et al. 2020)
@@
Bera et al. (2023) (a) Cross-center (4 Ep.) (b) cross-center (2 Ep.) (a) Binary prostate segmentation: Prostate158 (Adams et al. 2022) $\rightarrow$ NCI-ISBI (NCI ) $\rightarrow$ Promise12 (Litjens et al. 2014) $\rightarrow$ Decathlon (Antonelli et al. 2022) (b) binary hippocampus segmentation: Drayd (Denovellis et al. 2021) $\rightarrow$ HarP (Boccardi et al. 2015))
@@
Zhu et al. (2023) (a) Cross-site (6 Ep.) (b) cross-site & cross-modality (2 Ep.) (c) same-site & cross-modality (2 Ep.) (a) Binary prostate segmentation from T2-weighted MRI scans collected from 6 sites (12–30 scans/site) (Liu et al. 2020; Bloch et al. 2015; Lemaître et al. 2015) (b) multi-class (liver, left and right kidneys, and spleen) abdominal segmentation between 30 CT and 20 MRI T2-SPIR scans (c) muscle segmentation of 13 lower-leg muscles and bones between 30 MRI T1 and 30 mDixon scans
@@
Bándi et al. (2023) Cross-organ (3 Ep.) Histopathology data based cancer detection from breast (CAMELYON16, CAMELYON17), colon (private), and head-neck (private) datasets
@@ 3.4 Domain incremental scenarios
Domain Incremental Scenario (DIS) is the most popular and frequently observed category of CL scenarios for medical applications. This scenario aligns with the idea of learning in a changing environment where the datasets from different domains (e.g., research sites, hospitals, imaging modality, image acquisition protocol, etc.) are incrementally encountered over time, which thus involve covariance shift-induced discrepancies in data.
@@
Even if we have a single task and a fixed number of classes, these shifts can pose a challenge for the static deep model used in clinical settings because these shifts can quickly make existing models outdated and less effective. Here, a CL model aims to continuously update itself by incorporating new data streams that come from distributions that have shifted over time.
There have been plenty of attempts to develop CL models and evaluate different kinds of domain incremental scenarios. All the domain incremental scenarios are not equi-hard; some works consider a simpler level of domain shift, whereas some have a severe domain shift. To provide a comparative view, we tabulate all the domain incremental scenarios considered in literature through Table 5.
@@
Liu et al. (2023a) Brain tumor segmentation (3 Ep.) CIS with domain shift conditions: incrementally learn tumor core (BraTS2013 (Menze et al. 2014) dataset), the enhancing tumor (TCIA (Clark et al. 2013) dataset), and edema (CBICA (Bakas et al. 2018) dataset) structures via three datasets, each following different data distributions
Sadafi et al. (2023) WBC classification (3 Ep.) DIS (multi-site) + CIS using 3 WBC classification datasets having different number of classes: CIS on Matek-19 (Matek et al. 2019) then shift to INT-20 dataset and use it in CIS manner then shift to Acevedo-20 (Acevedo et al. 2020) dataset and use in CIS manner
Yang et al. (2023b) MedMNIST (Yang et al. 2023a) based disease classification (4 Ep.) CIS+DIS: 3 datasets (PathMNIST, DermaMNIST, and OrganAMNIST) as source domain and then 3 more datasets (RetinaMNIST, BreastMNIST, BloodMNIST) as domain shift but classes are added in class incremental fashion with 1 class at a time
Wang et al. (2023b) CIS (a) segmentation for endoscopy (2-3 Ep.) (b) surgical instrument segmentation (2 Ep.) Incrementally adding segmentation structure from EDD2020 (Ali et al. 2021, 2020) dataset which includes 5 classes as Barrett’s esophagus, cancer, high-grade dysplasia, polyp, and suspicious. Initially, 3-4 classes are in 1st episode, and then 2 or 1 classes are incrementally added in subsequent episodes. (b) EndoVis18 (Allan et al. 2020) and EndoVis17 (Allan et al. 2019) datasets as two episodes having some overlapping classes
Liu et al. (2023a) MRI-based brain tumor segmentation (4 Ep.) CIS+DIS Incrementally adding more structures to segmentation task from dataset acquired from different vendors/centers